# The genealogy of nearly critical branching processes in varying environment

### v3

@@ Abstract
Building on the spinal decomposition technique in [15] we prove a Yaglom limit law for the rescaled size of a nearly critical branching process in varying environment conditional on survival. In addition, our spinal approach allows us to prove convergence of the genealogical structure of the population at a fixed time horizon – where the sequence of trees are envisioned as a sequence of metric spaces – in the Gromov–Hausdorff–Prohorov (GHP) topology. We characterize the limiting metric space as a time-changed version of the Brownian coalescent point process [48]. Beyond our specific model, we derive several general results allowing to go from spinal decompositions to convergence of random trees in the GHP topology. As a direct application, we show how this type of convergence naturally condenses the limit of several interesting genealogical quantities: the population size, the time to the most-recent common ancestor, the reduced tree and the tree generated by $k$ uniformly sampled individuals. As in a recent article by the authors [15], we hope that our specific example illustrates a general methodology that could be applied to more complex branching processes.
@@ 1 Introduction
Describing the genealogies of different population models as the size of the population gets large is a prominent task in the theory of mathematical population genetics as well as in the study of branching processes, and goes back at least to the two seminal papers of Kingman [35, 36]. As of today there is a vast literature on the genealogies of various type of populations models, see for example [4, 23, 33, 46, 48, 49]. Here, we focus among other topics on the scaling limit of the genealogy of nearly critical branching processes in varying environment conditional on survival, for a precise definition see Section 2.
@@
The scope of this paper is two-fold. On the one hand we prove a Yaglom-type result [51] for the size of a nearly critical branching process in varying environment conditioned to survive for a long time, as well as a result on the scaling limit of the genealogy. For this purpose we prove a Kolmogorov-type asymptotics [37] for the probability of survival for an unusually long time, which might be interesting in its own right. On the other hand, our objective is to use this specific model to give an illustration of the techniques developed in [15, 19, 17, 22]. Our approach relies on new general results on random metric spaces that could be applied widely to branching processes with a critical behaviour. The spinal decomposition method in [15] gives an approach to study genealogies by showing convergence of random metric measure spaces in the Gromov-weak topology via a method of moments. To prove convergence of the $k$-th moment of a genealogy, a many-to-few formula allows us to reduce this problem to calculating the distribution of a tree with $k$ leaves, the so-called $k$-spine tree. The distribution of the $k$-spine tree is obtained through a random change of measure relying on choosing a suitable ansatz for the genealogy of the process. In order to make the $k$-spine method work one would require the offspring distribution to have moments of all order, which is far from the optimal conditions. However, using a truncation argument, we are able to work under a second moment condition. Finally, we show how to reinforce the Gromov-weak convergence of the genealogy to a Gromov–Hausdorff–Prohorov convergence. By doing so, we obtain as a direct consequence of our main result the convergence of many interesting genealogical quantities.
@@
In summary, a fraction of this paper will be dedicated to proving general results on random metric spaces. It is only at the end of the paper that this general formalism will be used to treat our specific problem as a (semi-)direct application. Along the way, we will obtain new criteria to obtain a Yaglom law for branching processes in varying environment (see Remark 2.2 below). We hope that this application will help to convince the reader that the abstract formalism introduced in [15, 19, 17, 22] (among others) and further developed in this paper is particularly well suited when treating general branching processes near criticality.
@@ 1.1 Relevant literature
The genealogical structure of branching processes is a topic that has drawn a lot of attention, especially in the critical case. Let us give a non-exhaustive account of some results related to our work. A first class of results focuses on the so-called reduced process [13, 52, 47], that counts the number of individuals having descendants at a fixed focal generation in the future. The limit of the reduced process is often expressed as a (time-inhomogeneous) pure birth process, see [33] for results in this direction for branching processes in varying environment. Second, in the same spirit as the theory of exchangeable coalescents, there has been an interest in describing the genealogy obtained by sampling a fixed number $k$ of individuals uniformly at a given time [29, 26, 40]. We also want to point at [38] for an explicit construction of the genealogy of splitting trees as a discrete coalescent point process, which plays a central role in [40].
@@
In this work, we show how the convergence of the previous two quantities, the reduced process and the genealogy of a uniform sample, as well as the convergence of the population size can be aggregated by viewing the population as a random ultrametric measure space converging in the Gromov–Hausdorff–Prohorov topology. Encoding trees as random metric spaces is a long-standing idea that has proved fruitful both for constant-size population models [11, 19] and for branching processes [2, 9]. Let us point at the following difference between our approach and more common convergence results in the Gromov–Hausdorff–Prohorov topology such as [2, 43, 45]. Here, we consider the genealogy of the population at a fixed time horizon, whereas many works [2, 43, 45] are interested in the limit of the whole genealogical tree. Although in the limit one can construct the genealogy at a fixed time (the Brownian coalescent point process) from that of the whole tree (the Brownian continuum random tree), see [48], it is not straightforward to go from the convergence of the latter object to our type of convergence result. Also, from a technical point of view, the convergence of the whole tree is typically obtained by studying a height function, whereas our approach relies on a $k$-spine method.
@@
The spinal decomposition of branching processes is widely used in the theory of branching processes. In particular, the case of a single spine has been proven to be an important tool, see [44, 50, 16]. The classical Yaglom limit and the Kolmogorov estimate for critical Galton–Watson processes is obtained in [44] using this technique. Here, we adapt the one spine decomposition from [44] to prove the Kolmogorov asymptotics for branching processes in varying environment, see Theorem 1. These techniques have been developed further for example in [27, 15] to allow for a $k$-spine construction, which in turn allows us to compute higher moments of the tree by a random change of measure. Such a $k$-spine construction has also been used to extract information about the genealogy of several branching processes [26, 29, 15], and the current work follows a similar path. We also want to point at [18, 25] for a computation of higher order moments for general spatial branching processes, and the derivation of a Yaglom law in this context.
@@
For an introduction to branching processes in varying environment (BPVE) we refer to the monograph of Kersting and Vatutin [34]. The work of Kersting [32] gives a complete classification of critical branching processes in varying environment (among the other cases, supercritical, subcritical and asymptotically degenerate) and proves a Yaglom limit result in this case. Later, Cardona-Tobón and Palau in [7] managed to prove the Yaglom-type result for critical branching processes in varying environment by applying a two-spine decomposition and therefore providing a probabilistic proof of the result obtained by analytical methods in [32]. They make use of the remarkable fact that the distribution of the time to the most recent common ancestor of two individuals can be found explicitly. Recently, in [33] Kersting managed to obtain the asymptotic distribution of the time to the most recent common ancestor of all individuals living at some large time $n$. Our work is a continuation and extension of these results, but does not build strongly on them, since we are dealing with nearly critical processes instead.
@@
In the process of finishing this article we became aware of the works of Harris, Palau and Pardo [24] and of Conchon–Kerjan, Kious and Mailler [8]. In the former work, building on techniques of theirs developed in [27] and [7], the authors give a forward in time construction of the $k$-spine for a critical branching process in varying environment. Subsequently, via a change of measure relating the original branching process in varying environment and the $k$-spine tree they obtain their main result on the asymptotic distribution describing the splitting times for a uniform $k$-sample from the tree at large times, compare Corollary 2.3 (iii). Their results and the spinal decomposition are related to our work and are in the same spirit. However, due to the different points of view and approaches in their work and ours, we believe that these two papers can be seen as complementary to each other. We refer to Remark 2.2 and Remark 2.4 for a more in depth comparison. In the latter work [8], the authors derive the scaling limit of the whole tree structure of a branching process in i.i.d. random environment for the Gromov–Hausdorff–Prohorov topology. Their main result is consistent with ours in the sense that, under their hypothesis, the limit that we find (the Brownian coalescent point process) corresponds to the reduced tree at a given time of their limit (the Brownian continuum random tree). Nevertheless, as mentioned above, our type of result is not directly implied by the convergence of the whole tree and the techniques are different. Also, our hypothesis can be seen as more general, in the sense that we do not require the environment to have properties of i.i.d. sequences, and consider a sequence of nearly critical environments rather than a single strictly critical environment. In particular, stationary environments (and thus i.i.d. environments) can only lead to the Brownian coalescent point process in the limit, whereas we recover any time-change of a Brownian coalescent point process, including non-binary trees whenever the limiting variance process has jumps. Finally, let us point at [20] which considers a binary branching process in a random environment, modelling the effect of a catalyst that modifies the branching rate of the population. They provide the scaling limit of the genealogy of this branching process, recovering in the limit a time-changed version of a coalescent point process similar to the one arising here. Our work goes way beyond the case of binary branching processes covered in [20], and in a sense extends this result to a much broader universality class.
@@ 1.2 Outline
The outline of this paper is as follows. First, in Section 2 we will define precisely the notion of a nearly critical branching process in varying environment and give some examples. We also state our two main results, the Kolmogorov asymptotic (Theorem 1) and the convergence of the genealogy (Theorem 2), and give in Corollary 2.3 straightforward consequences of our Gromov–Hausdorff–Prohorov convergence. In Section 3 we introduce the topologies on metric measure spaces needed in this work, and prove some continuity results for ultrametric spaces. The limiting object for the genealogy, the environmental coalescent point process, is introduced in Section 4, where we also derive some of its properties. The last sections are devoted to the proofs of our results. In Section 5 we recall the spinal decomposition tools for branching processes that we need. The proof of the Kolmogorov asymptotics (Theorem 1) is carried out in Section 6, and that of the convergence of the genealogy (Theorem 2) in Section 7. Lastly, some results of technical nature are given in the Appendix A.
@@ 2 Model
Branching processes in varying environment (BPVE for short) are a natural generalisation of classical Galton–Watson processes, which allow for a changing offspring distribution between the generations. Throughout this work we will try to follow the notation of [34].
Let us start by properly defining the notion of a branching process in varying environment. Denote by $\mathcal{P}(\mathbb{N}_{0})$ the space of probability measures on $\mathbb{N}_{0}=\{0,1,2,\dots\}$ and let $v=(f_{1},f_{2},\dots)$ be a sequence of probability measures on $\mathbb{N}_{0}$.
@@
where for each $n\geq 1$ the family $\{\xi_{i,n},i\geq 1\}$ consists of independent random variables with distribution $f_{n}$.
Note that we enforce the BPVE to consist of a single individual in generation 00. In the subsequent parts it will be useful to identify measures $f\in\mathcal{P}(\mathbb{N}_{0})$ with their generating function
@@
where $f[k]$, $k\geq 0$, denotes the weight of the distribution $f$ in $k$. Note that we will use the same symbol $f$ for the distribution as well as for the corresponding generating function $f(s)$. Recall that the first and second factorial moments of a random variable with distribution $f$ can be expressed as
@@
In the following we will consider a sequence $Z^{(N)}=\big{(}Z_{n}^{(N)},n\geq 0\big{)}$ of branching processes in varying environment indexed by $N$. That is, we consider a sequence of environments $v^{(N)}=(f_{1}^{(N)},f_{2}^{(N)},\dots)$ and study the asymptotic probability for this process to survive $N$ generations as well as the asymptotic genealogy of the population at a large time.
Due to the fact that we are working with a sequence of branching processes, all the quantities introduced so far as well as the quantities which will be introduced in the subsequent sections depend on $N$. Additionally, all expectations and probabilities will depend on $N$. For the convenience of the reader and to avoid overloading the notation, we will drop the superscript $N$ most of the times. The reader should keep the implicit dependency of $N$ in mind. Furthermore, whenever there is a sum ranging from $1$ to $\lfloor Nt\rfloor$ we will abuse the notation and simply write $Nt$ as the upper limit. More generally, whenever no confusion is possible, we simply write $Nt$ instead of $\lfloor Nt\rfloor$, as in $f_{Nt}$ or $\xi_{Nt}$.
@@
We can construct a random tree out of an independent collection $(K_{u},u\in\mathcal{U})$ of r.v. with $K_{u}\sim f_{|u|+1}$ by setting
$T=\{u=(u_{1},\dots,u_{n}):\forall i\leq n,u_{i}\leq K_{(u_{1},\dots,u_{i-1})}\}.$ (4)
@@ 2.1 Assumptions
In order to prove the main theorems, Theorem 1 and Theorem 2, we will work under the following assumptions on the sequence of environments $v=(f_{1},f_{2},\dots)$, which we name a nearly critical varying environment. Recall that we do not denote the dependency of $v$ on $N$ explicitly. Let us introduce the following quantities,
@@
which is the expectation of the BPVE at time $k$, precisely $\mathbb{E}\left[Z_{k}\right]=\mu_{k}$. We call a sequence of branching processes in varying environment nearly critical, if
@@
in the Skorohod topology for some càdlàg process $(X_{t},t\geq 0)$. (Having in mind the important case where the environment $v$ is a realization of an i.i.d. sequence, we use the terminology process for $(X_{t},t\geq 0)$ although it is a deterministic càdlàg function). The next assumptions ensure that the variance between the generations does not fluctuate too strongly. Precisely, for some càdlàg non-decreasing function $\sigma^{2}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ and some sequence $\kappa_{N}\to\infty$, we assume that for all $t\in\mathbb{R}_{+}$ such that $\sigma^{2}$ is continuous at $t$,
@@
We also assume that $\sigma^{2}$ and $(X_{t},t\geq 0)$ do not jump at the same time. Lastly, we will enforce the following Lindeberg-type condition
@@
which prevents any single individual from having a large number of offspring of order $\kappa_{N}$. We believe that given (6) and (7) this assumption, which was already used in [6], is nearly optimal. For instance, for Galton–Watson processes (8) is known to be optimal for the process started from a population size of order $\kappa_{N}=N$ to converge to a Feller diffusion [21].
@@
Our definition of near criticality is directly inspired by that for Galton–Watson processes, where one generally imposes that the mean and the variance of the process converge. An alternative notion of criticality was proposed for BPVEs in [32] and was also further used in [33, 7, 24]. It relies on the behaviour of the following two series
@@
where the asymptotics under assumptions (6) and (7) are obtained in Lemma A.4. Since both series diverge, our BPVE would also be classified as critical in [32]. Assumptions (6) and (7) are not necessary for these series to diverge and can thus be thought of as restrictive in that sense. It should be noted however that these stronger assumptions allow us to work with sequences of environments and lead to a notion of near criticality, whereas the aforementioned works consider fixed environment. Also, we only need the mild moment condition (8) on the offspring size which we believe to be nearly optimal, whereas [7, 24] require a finite third moment, and [32, 33] a stronger second moment assumption. In particular, if (6) and (7) hold, the latter condition (assumption (B) in [32]) implies that for any sequence $K_{N}\to\infty$
@@
see Lemma A.6. Compare this to the weaker assumptions (8) and (8). In particular, (9) prevents the limit process $\sigma^{2}$ from jumping, which in turn implies that the limiting tree is binary.

### v4

@@ Abstract
Building on the spinal decomposition technique in [15] we prove a Yaglom limit law for the rescaled size of a nearly critical branching process in varying environment conditional on survival. In addition, our spinal approach allows us to prove convergence of the genealogical structure of the population at a fixed time horizon – when the sequence of trees are envisioned as a sequence of metric spaces – in the Gromov–Hausdorff–Prohorov (GHP) topology. We characterize the limiting metric space as a time-changed version of the Brownian coalescent point process [48].
Beyond our specific model, we derive several general results allowing one to go from spinal decompositions to convergence of random trees in the GHP topology. As a direct application, we show how this type of convergence naturally condenses the limit of several interesting genealogical quantities: the population size, the time to the most-recent common ancestor, the reduced tree, and the tree generated by $k$ uniformly sampled individuals. As in a recent article by the authors [15], we hope that our specific example illustrates a general methodology that could be applied to more complex branching processes.
@@ 1 Introduction
Describing the genealogies of different population models as the size of the population gets large is a prominent task in the theory of mathematical population genetics as well as in the study of branching processes, and goes back at least to the two seminal papers of Kingman [35, 36]. As of today, there is a vast literature on the genealogies of various types of population models, see for example [4, 23, 33, 46, 48, 49]. Here, among other topics, we focus on the scaling limit of the genealogy of nearly critical branching processes in varying environment conditional on survival, see Section 2 for a precise definition.
@@
The scope of this paper is two-fold. On the one hand, we prove a Yaglom-type result [51] for the size of a class of nearly critical branching processes in varying environment conditioned on survival for a long time, as well as a result on the scaling limit of their genealogy. For this purpose we prove a Kolmogorov-type asymptotics [37] for the probability that the process survives for a long time, which might be interesting in its own right. On the other hand, our objective is to use this specific model to illustrate the techniques developed in [15, 19, 17, 22]. Our approach relies on new general results on random metric spaces that could be applied widely to branching processes with critical behavior. The spinal decomposition method in [15] enables one to study genealogies by showing convergence of random metric measure spaces in the Gromov-weak topology using a method of moments. To prove convergence of the $k$-th moment of a genealogy, a many-to-few formula allows us to reduce this problem to calculating the distribution of a tree with $k$ leaves, the so-called $k$-spine tree. The distribution of this $k$-spine tree is obtained through a random change of measure relying on choosing a suitable ansatz for the genealogy of the process. In order to make the $k$-spine method work one would require the offspring distribution to have moments of all order, which is far from the optimal condition. However, using a truncation argument, we are able to work under a second moment condition. Finally, we show how to reinforce the Gromov-weak convergence of the genealogy to a Gromov–Hausdorff–Prohorov convergence. As a direct consequence of our main result, we obtain the convergence of many interesting genealogical quantities.
@@
In summary, a part of the current work is dedicated to proving general results on random metric spaces. It is only at the end of the article that this general formalism will be used to treat our specific problem as a (semi-)direct application. Along the way, we will obtain new criteria to obtain a Yaglom law for branching processes in varying environment (see Remark 2.2 below). We hope that this application will help convince the reader that the abstract formalism introduced in [15, 19, 17, 22] (among others) and further developed here is particularly well suited when treating general branching processes near criticality.
@@ 1.1 Relevant literature
The genealogical structure of branching processes is a topic that has drawn considerable attention, particularly in the critical case. We give a non-exhaustive overview of some results related to our work. A first class of results focuses on the so-called reduced process [13, 52, 47], that counts the number of individuals having descendants at a fixed generation in the future. The limit of the reduced process is often expressed as a (time-inhomogeneous) pure birth process, see [33] for results in this direction for branching processes in varying environment. Second, in the same spirit as the theory of exchangeable coalescents, there has been an interest in describing the genealogy obtained by sampling a fixed number $k$ of individuals uniformly at a given time [29, 25, 40]. We also want to point at [38] for an explicit construction of the genealogy of splitting trees as a discrete coalescent point process, which plays a central role in [40].
@@
In this work, we show how the convergence of the reduced process, of the genealogy of a uniform sample, and of the population size can be deduced by viewing the population as a random ultrametric measure space converging in the Gromov–Hausdorff–Prohorov topology. Encoding trees as random metric spaces is a long-standing idea that has proved fruitful both for constant-size population models [11, 19] and for branching processes [2, 9]. Let us point out the following difference between our approach and more common convergence results in the Gromov–Hausdorff–Prohorov topology such as [2, 43, 45]. Here, we consider the genealogy of the population at a fixed time horizon, whereas many works [2, 43, 45] are interested in the limit of the whole genealogical tree. Although in the limit one can construct the genealogy at a fixed time (the Brownian coalescent point process) from that of the whole tree (the Brownian continuum random tree), see [48], it is not straightforward to go from the convergence of the latter object to our type of convergence result. Also, from a technical point of view, the convergence of the whole tree is typically obtained by studying a height function, whereas our approach relies on a $k$-spine method.
@@
Spinal decompositions techniques are widely used in the theory of branching processes. In particular, the case of a single spine has been proven to be an important tool, see [44, 50, 16]. The classical Yaglom limit and the Kolmogorov estimate for critical Galton–Watson processes are obtained in [44] using this technique. Here, we adapt the one-spine decomposition from [44] to prove the Kolmogorov asymptotics for branching processes in varying environment, see Theorem 1. These techniques have been further developed – for example in [27, 15] – to allow for a $k$-spine construction, which in turn allows us to compute higher moments of the tree by a random change of measure. Such a $k$-spine construction has also been used to extract information about the genealogy of several branching processes [25, 29, 15], and the current work follows a similar path. We also want to point at [18, 24] for a computation of higher order moments for general spatial branching processes, and the derivation of a Yaglom law in this context.
@@
For an introduction to branching processes in varying environment we refer to the monograph of Kersting and Vatutin [34]. The work of Kersting [32] gives a complete classification of critical branching processes in varying environment and proves a Yaglom limit result in this case. Later, Cardona-Tobón and Palau [7] managed to prove the Yaglom-type result for critical branching processes in varying environment by applying a two-spine decomposition and therefore providing a probabilistic proof of the result obtained by analytical methods in [32]. They make use of the remarkable fact that the distribution of the time to the most-recent common ancestor of two individuals can be found explicitly. Recently, Kersting [33] managed to obtain the asymptotic distribution of the time to the most-recent common ancestor of all individuals living at some large generation $n$. Our work is a continuation and extension of these results, but does not build strongly on them, since we are dealing with nearly critical processes instead.
@@
While completing this article we became aware of the works of Harris, Palau and Pardo [26] and of Conchon–Kerjan, Kious and Mailler [8]. In the former work, building on techniques of theirs developed in [27] and [7], the authors give a forward in time construction of the $k$-spine for a critical branching process in varying environment. Subsequently, via a change of measure relating the original branching process in varying environment and the $k$-spine tree, they obtain the asymptotic distribution of the splitting times of the tree spanned by $k$ individuals chosen uniformly at large times, which bares similarities to our Corollary 2.3 (iii). Their results and the spinal decomposition are related to our work and are in the same spirit. However, due to the different points of view and approaches in their work and ours, we believe that these two articles are complementary to each other. We refer to Remark 2.2 and Remark 2.4 for a more detailed comparison. In the latter work [8], the authors derive the scaling limit of the whole tree structure of a branching process in i.i.d. random environment for the Gromov–Hausdorff–Prohorov topology. Their main result is consistent with ours in the sense that, under their hypothesis, the limit that we find (the Brownian coalescent point process) corresponds to the reduced tree at a given time of their limit (the Brownian continuum random tree). Nevertheless, as mentioned above, our type of result is not directly implied by the convergence of the whole tree and the techniques are different. Also, our hypothesis can be seen as more general, in the sense that we do not require the environment to have properties of i.i.d. sequences, and consider a sequence of nearly critical environments rather than a single strictly critical environment. In particular, stationary environments (and thus i.i.d. environments) can only lead to the Brownian coalescent point process in the limit, whereas we recover any time-change of a Brownian coalescent point process, including non-binary trees whenever the limiting variance process has jumps. Finally, let us point to [20] which considers a binary branching process in a random environment, modelling the effect of a catalyst that modifies the branching rate of the population. They provide the scaling limit of the genealogy of this branching process, recovering in the limit a time-changed version of a coalescent point process similar to the one arising here. Our work goes way beyond the case of binary branching processes covered in [20], and in a sense extends this result to a much broader universality class.
@@ 1.2 Outline
The outline of the manuscript is as follows. First, in Section 2 we precisely define the notion of a nearly critical branching process in varying environment and provide some examples. We also state our two main results, the Kolmogorov asymptotics (Theorem 1) and the convergence of the genealogy (Theorem 2), and in Corollary 2.3 we state some straightforward consequences of our Gromov–Hausdorff–Prohorov convergence. In Section 3 we introduce the topologies on metric measure spaces needed in this work, and prove some continuity results for ultrametric spaces. In Section 4, we introduce the limiting object for the genealogy, the environmental coalescent point process and derive some of its properties. The final sections are devoted to the proofs of our results. In Section 5 we recall the spinal decomposition tools for branching processes that we need. The proof of the Kolmogorov asymptotics (Theorem 1) is carried out in Section 6, and that of the convergence of the genealogy (Theorem 2) in Section 7. Lastly, some results of technical nature are given in Appendix A.
@@ 2 Model
Branching processes in varying environment (BPVE for short) are a natural generalization of classical Galton–Watson processes in which the offspring distribution can change between generations. Throughout this work we will try to follow the notation of [34].
We begin by defining formally the notion of a branching process in varying environment. Denote by $\mathcal{P}(\mathbb{N}_{0})$ the space of probability measures on $\mathbb{N}_{0}=\{0,1,2,\dots\}$ and let $v=(f_{1},f_{2},\dots)$ be a sequence of probability measures on $\mathbb{N}_{0}$.
@@
where for each $n\geq 1$ the random variables $\{\xi_{i,n},i\geq 1\}$ are independent and distributed as $f_{n}$.
Note that we assume that the BPVE starts from a single individual at generation 0. Subsequently, it will be useful to identify a measure $f\in\mathcal{P}(\mathbb{N}_{0})$ with its generating function
@@
where $f[k]$, $k\geq 0$, denotes the weight of the distribution $f$ at $k$. Note that we will use the same symbol $f$ for the distribution as well as for the corresponding generating function $f(s)$. Recall that the first and second factorial moments of a random variable with distribution $f$ can be expressed as
@@
In the following, we consider a sequence $Z^{(N)}=\big(Z_{n}^{(N)},n\geq 0\big)$ of branching processes in varying environment indexed by $N$. That is, we consider a sequence of environments $v^{(N)}=(f_{1}^{(N)},f_{2}^{(N)},\dots)$ and study the asymptotic probability that this process survives $N$ generations as well as the asymptotic genealogy of the population at large times.
Due to the fact that we are working with a sequence of branching processes, all the quantities introduced so far as well as the quantities which will be introduced in the subsequent sections depend on $N$. Additionally, all expectations and probabilities will depend on $N$. For the convenience of the reader and to avoid overloading the notation, we will drop the superscript $N$ most of the times. The reader should keep the implicit dependency on $N$ in mind. Furthermore, whenever a sum is ranging from $1$ to $\lfloor Nt\rfloor$, we will abuse the notation and simply write $Nt$ as the upper limit. More generally, whenever no confusion is possible, we simply write $Nt$ instead of $\lfloor Nt\rfloor$, as in $f_{Nt}$ or $\xi_{Nt}$.
@@
We can construct a random tree out of an independent collection $(K_{u},u\in\mathcal{U})$ of random variables with $K_{u}\sim f_{|u|+1}$ by setting
$T=\{u=(u_{1},\dots,u_{n}):u_{i}\leq K_{(u_{1},\dots,u_{i-1})},\>i\leq n\}.$ (4)
@@ 2.1 Assumptions
In order to prove the main theorems, Theorem 1 and Theorem 2, we will work under the following assumptions on the sequence of environments $v=(f_{1},f_{2},\dots)$, which we call a nearly critical varying environment. Recall that we do not denote the dependency of $v$ on $N$ explicitly. Let us introduce the following quantities,
@@
which is the expectation of the BPVE at time $k$, namely $\mathbb{E}\left[Z_{k}\right]=\mu_{k}$. We call a sequence of branching processes in varying environment nearly critical if
@@
in the Skorohod topology for some càdlàg process $(X_{t},t\geq 0)$. (Having in mind the important case where the environment $v$ is a realization of an i.i.d. sequence, we use the terminology process for $(X_{t},t\geq 0)$ although it is a deterministic càdlàg function). The following assumptions ensure that the variance between the generations does not fluctuate too strongly. Precisely, for some càdlàg non-decreasing function $\sigma^{2}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ and some sequence $\kappa_{N}\to\infty$, we assume that for all $t\in\mathbb{R}_{+}$ such that $\sigma^{2}$ is continuous at $t$,
@@
We also assume that $\sigma^{2}$ and $(X_{t},t\geq 0)$ do not jump at the same time. Lastly, we impose the following Lindeberg-type condition
@@
which prevents any single individual from having a large number of offspring of order $\kappa_{N}$. We believe that given (6) and (7) this assumption, already used in [6], is nearly optimal. For instance, in the case of Galton–Watson processes (8) is known to be optimal for convergence to a Feller diffusion when the initial population size is of order $\kappa_{N}$ [21].
@@
Our definition of near criticality is directly inspired by that for Galton–Watson processes, where one generally imposes that the mean and the variance of the process converge. An alternative notion of criticality was proposed for BPVEs in [32] and was further used in [33, 7, 26]. It relies on the behavior of the following two series
@@
where the asymptotics under assumptions (6) and (7) are obtained in Lemma A.4. Since both series diverge, our BPVE would be classified as critical in [32]. Assumptions (6) and (7) are not necessary for these series to diverge and can therefore be seen as restrictive in that sense.
It should be noted, however, that these stronger assumptions allow us to work with sequences of environments and lead to a notion of near criticality, whereas the aforementioned works consider a fixed environment. Moreover, we only require the mild moment condition (8) on the offspring size, which we believe to be nearly optimal, whereas [7, 26] require a finite third moment, and [32, 33] a stronger second moment assumption. In particular, if (6) and (7) hold, the latter condition (assumption (B) in [32]) implies that for any sequence $K_{N}\to\infty$
@@
see Lemma A.6. Compare this with the weaker assumptions (8) and (8). In particular, (9) prevents the limiting process $\sigma^{2}$ from jumping, which in turn implies that the limiting tree is binary.